OpenStudy (anonymous):

Find the ratio of successive terms. Write your answer as a fully simplified fraction. the ratio successive term for 2^(n)/(2n)! is (2)/(2n+1) RIGHT?

OpenStudy (anonymous):

No, that's not correct, $$ \large \frac{2^n} {2n!} =\frac{2^n} {1.3.5.7 \cdots (2n-1) \times 2.4.6.8. \cdots 2n} $$ factor out 2^n from both numerator and denominator.. you will have you answer.

jhonyy9 (jhonyy9):

1. 2/2 =1 2. 4/(2(1*2))=4/4=1 3. 8/(2(1*2*3))=8/12 =2/3 4 .......